# Extended Reality for Enhanced Human-Robot Collaboration: a Human-in-the-Loop Approach

### v1

@@ Abstract
I INTRODUCTION
The transition from Industry 4.0, which has facilitated the development and implementation of autonomous cyber-physical systems, IoT, and Big Data in manufacturing, to Industry 5.0 that aims to complement the technological advancements by prioritizing human-centric approaches, fundamentally reshaping the interaction between humans and machines within the manufacturing sector. This evolution involves combining the precision of robots and machines with the intelligence and versatility of human input [1]. In this case, one of the most important challenges is the design of communication interfaces that accurately represent the manufacturing processes, account for needs in adaptability and flexibility, and provide intuitive interaction methods for the users. The primary focus of most robot software lies in programming specific functions, with limited to no tolerance for deviations from the programmed settings. Machine learning (ML) offers new possibilities by enhancing the generalization capabilities of robots’ decision-making. Nevertheless, it is crucial to recognize ML as an enabling tool rather than as the sole medium for human-robot communication.
@@ I INTRODUCTION
The complexity of HRC scenarios arises from the diverse range of tasks, environments, and the varying capabilities that both humans and robots bring to collaborative efforts. Multiple studies have been conducted in an attempt to define the levels of interaction in human-robot collaboration processes [6, 7]. However, the majority of existing research predominantly addresses specific facets of human-robot interaction (HRI), potentially overlooking the broader span of interaction dynamics. A more flexible approach focusing on industrial settings is proposed by Mukherjee et al. in [8], where authors’ classification criteria include the characteristics of the task and the workspace, autonomy level and operational mode of the robot, and the allowance for physical contact between the agents. This taxonomy ranges from level zero, being fully programmed robotics, to level five, classified as completely autonomous robots. Interestingly, both ends of the spectrum envisage no human interaction. Level four, the collaboration, involves humans and robots working simultaneously towards common goals. A more detailed and organized view of the approach from [8] is shown in Table I.
TABLE I: Industrial HRI levels, adopted from [8]
Level
Interaction
Description
L0
Fully Programmed
Traditional approach with physically restricting cages, no consideration of HRC.
L1
Co-existence
The agents are separated by safety zone, the robot pauses its operation in case human enters the area.
L2
Assistance
Robot can assist human in a certain task (such as operations with heavy objects), but it has no independent tasks.
L3
Co-operation
The agents work together towards the common goal within the designated intervention zone. However, human and robot do not share the same task and there is no physical contact.
L4
Collaboration
Humans and robots working autonomously towards the same goal sharing the the task, workspace, and resources.
L5
Fully Autonomous
Generalizable manipulators trained using ML algorithms, no human intervention is considered.
Moreover, we believe that the synergy between machine learning and extended reality presents a unique potential to provide an intuitive approach for human operators to act as robots’ instructors. In this work, we propose an additional extension of the above-mentioned levels from [8], four and five, named ”fully autonomous with the human-in-the-loop.” While this sub-level may not require a continuous human-robot interaction, it does offer the possibility for a human operator to step in during the autonomous process, augmenting it with human expertise as needed.
@@
II FRAMEWORK CONCEPTUALIZATION
Defining autonomous robotics is a critical aspect of this paper.Hence, Section II-A is dedicated to introducing the machine learning-based manipulator, and the respective framework for tasks generalization. Simultaneously, Section II-B delves into the integration of the human component through XR, specifically focusing on collaboration, programming, and performance oversight.
II-A Manipulator Task Generalization
@@ II-A1 Demo
II-A2 Task Planner
It is assumed that the actual task executes similar skills compared to the demo task, but in a different environment (e.g., varied toolsets, obstacles, and parametric characteristics). Therefore, the Task Planner decomposes the provided demonstration (i.e., sample task) into multiple skills, which are then associated with and utilized for the execution of the actual task. Usually, skills are defined as low-level abstractions or primitives [12]; for example, the task is decomposed into the commands to “move linearly”, “locate the object”, “position the gripper”, “pick-up the object”, etc. Hence, the skills library is created, as shown in the Fig. 1.
II-A3 Task Execution
@@
II-B Human-in-the-Loop Component
The use of immersive technology that augments or completely replaces the real world opens up opportunities for safe and intuitive human-robot interaction and task programming. We present XR-based interaction approaches with the goal of integrating human intelligence into the robot’s learning process for task generalization. Also, the methods described below are easily translated to various pre-trained autonomous manipulators. The heightened control allows for easy customization to meet task requirements in a safer and more adaptive manner. We further elaborate on the conceptual basis for each method outlined in Fig. 1.
II-B1 Immersive Demonstration
@@ III APPLICATIONS REVIEW
III-A Operator Support and Communication
@@ III-B Safety Considerations
Some authors approach the concerns for safety from another corner of the XR paradigm - Virtual Reality. Creating a completely virtual environment allows the mimicking of realistic as well as potential hypothetical scenarios [7], where users can interact and familiarize themselves with the equipment at no risk of injury. Additionally, the concept of HRI in virtual reality can be elevated by incorporating the digital twin of the manipulator. Hence, it is no longer just a training simulation, but a real-time teleoperation (discussed in detail in III-C) that blurs the boundary between the virtual and physical interaction [29]. On the other hand, the aspects of mental and physical load of interactions in VR are not fully examined, and concerns of cybersickness should be addressed through further research [30].
III-C Teleoperation
@@ IV DISCUSSION
IV-A Mitigating Risks
The visualization of motion intention and object manipulation gives the operator a better understanding of the workflow. In the case of VR systems, one of the main advantages is the elimination of the need for physical expert presence [45]. In general, studies show that XR is a unique tool that allows the conduct of teleoperation, robot programming, and various operator-supporting functions in a safe and controlled manner. Although XR offers the possibility of finding a balance between operator safety and robot efficiency, the impact on physical and mental health from working with head-mounted displays in HRI tasks necessitates further studies. As mentioned in [30], there are currently no optimal solutions to address all possible side effects like muscle fatigue, motion sickness, and mental overload. Similarly, for human-robot collaborative tasks, the psychological factor remains a significant area of research. This includes the effort to cultivate trust between the agents, as well as methods to generate motion and trajectories that closely mimic human behavior.
IV-B Immersiveness
One of the bases for introducing XR is its immersive potential. The immersive environment lets us maintain control over all simulation parameters, which is particularly important in the context of immersive demonstrations. As it has been shown in the current literature [38, 46], demonstration quality is one of the most decisive factors in the effectiveness of robot policy learning.
Considering enhanced visualization and situation awareness provided by the immersive capabilities, task coordination is another area of research for improving collaboration between the agents. Further investigation would look into the application of XR in the context of multi-robot systems. The ability to simulate and coordinate multi-robot systems would allow the integration of human operators into more complex manufacturing environments.
IV-C User-oriented Concepts
@@
IV-D From Lab to Industry
One of the primary obstacles preventing the adaptation of extended reality for the human-robot collaboration scenarios is setup costs, requiring substantial initial investment. Additionally, there is a lack of unified consensus within industry on the XR integration strategy and interface development, which understandably stems from the differences in industry-specific requirements and individual products. A possible future area of research could involve investigating the feasibility of utilizing the same XR interface across multiple manipulators with the flexibility to easily customize the interface for new robot- or product-specific characteristics. Furthermore, the implementation of XR is usually performed in a bundle with other digital technologies, as outlined in II-C. Therefore, it entails additional investments and resources, and in fact, the effectiveness of XR becomes conditional on the development and maintenance of other technologies. Also, it is noteworthy that most of the reviewed works perform experiments in the laboratory. The prospect of transferring the developed methodologies and interfaces to real-world scenarios and industrial settings remains yet to be explored.
V CONCLUSIONS

### v2

@@ Abstract
©2024 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
1 INTRODUCTION
The transition from Industry 4.0, which has facilitated the development and implementation of autonomous cyber-physical systems, IoT, and Big Data in manufacturing, to Industry 5.0 which aims to complement the technological advancements by prioritizing human-centric approaches, fundamentally reshapes the interaction between humans and machines within the manufacturing sector. This evolution involves combining the precision of robots and machines with the intelligence and versatility of human input [1]. In this case, one of the most important challenges is the design of communication interfaces that accurately represent the manufacturing processes, account for needs in adaptability and flexibility, and provide intuitive interaction methods for the users. The primary focus of most robot software lies in programming specific functions, with limited to no tolerance for deviations from the programmed settings. Machine learning (ML) offers new possibilities by enhancing the generalization capabilities of robots’ decision-making. Nevertheless, it is crucial to recognize ML as an enabling tool rather than as the sole medium for human-robot communication.
@@ 1 INTRODUCTION
The complexity of HRC scenarios arises from the diverse range of tasks, environments, and the varying capabilities that both humans and robots bring to collaborative efforts. Multiple studies have been conducted in an attempt to define the levels of interaction in human-robot collaboration processes [6, 7]. However, the majority of existing research predominantly addresses specific facets of human-robot interaction (HRI), potentially overlooking the broader span of interaction dynamics. A more flexible approach focusing on industrial settings is proposed by Mukherjee et al. in [8], where authors’ classification criteria include the characteristics of the task and the workspace, autonomy level and operational mode of the robot, and the allowance for physical contact between the agents. This taxonomy ranges from level zero, being fully programmed robotics, to level five, classified as completely autonomous robots. Interestingly, both ends of the spectrum envisage no human interaction. Level four, the collaboration, involves humans and robots working simultaneously towards common goals. A more detailed and organized view of the approach from [8] is shown in Table 1.
Table 1: Industrial HRI levels, adopted from [8]
Level Interaction Description
L0 Fully Programmed Traditional approach with physically restricting cages, no consideration of HRC.
L1 Co-existence The agents are separated by safety zone, the robot pauses its operation in case human enters the area.
L2 Assistance Robot can assist human in a certain task (such as operations with heavy objects), but it has no independent tasks.
L3 Co-operation The agents work together towards the common goal within the designated intervention zone. However, human and robot do not share the same task and there is no physical contact.
L4 Collaboration Humans and robots working autonomously towards the same goal sharing the the task, workspace, and resources.
L5 Fully Autonomous Generalizable manipulators trained using ML algorithms, no human intervention is considered.
Moreover, we believe that the synergy between machine learning and extended reality presents a unique potential to provide an intuitive approach for human operators to act as robots’ instructors. In this work, we propose an additional extension of the above-mentioned levels from [8], four and five, termed "fully autonomous with the human-in-the-loop." While this sub-level may not require a continuous human-robot interaction, it does offer the possibility for a human operator to step in during the autonomous process, augmenting it with human expertise as needed.
@@
2 FRAMEWORK CONCEPTUALIZATION
Defining autonomous robotics is a critical aspect of this paper. Hence, Section 2.1 is dedicated to introducing the machine learning-based manipulator, and the respective framework for tasks generalization. Simultaneously, Section 2.2 delves into the integration of the human component through XR, specifically focusing on collaboration, programming, and performance oversight.
2.1 Manipulator Task Generalization
@@ 2.1.1 Demo
2.1.2 Task Planner
It is assumed that the actual task executes similar skills compared to the demo task, but in a different environment (e.g., varied toolsets, obstacles, and parametric characteristics). Therefore, the Task Planner decomposes the provided demonstration (i.e., sample task) into multiple skills, which are then associated with and utilized for the execution of the actual task. Usually, skills are defined as low-level abstractions or primitives [12]; for example, the task is decomposed into the commands to “move linearly”, “locate the object”, “position the gripper”, “pick up the object”, etc. This decomposition forms the foundation of the skills library, a collection of fundamental actions essential for task completion. By providing crucial information, the skills library enhances the efficiency of task execution in different environments.
2.1.3 Task Execution
@@
2.2 Human-in-the-Loop Component
The use of immersive technology that augments or completely replaces the real world opens up opportunities for safe and intuitive human-robot interaction and task programming. We present XR-based interaction approaches with the goal of integrating human intelligence into the robot’s learning process for task generalization. Also, the methods described below are easily translated to various pre-trained autonomous manipulators. We further elaborate on the conceptual basis for each method outlined in Fig. 1.
2.2.1 Immersive Demonstration
@@ 3 APPLICATIONS REVIEW
3.1 Operator Support and Communication
@@ 3.2 Safety Considerations
Some authors approach the concerns for safety from another corner of the XR paradigm - Virtual Reality. Creating a completely virtual environment allows the mimicking of realistic as well as potential hypothetical scenarios [7], where users can interact and familiarize themselves with the equipment at no risk of injury. Additionally, the concept of HRI in virtual reality can be elevated by incorporating the digital twin of the manipulator. Hence, it is no longer just a training simulation, but a real-time teleoperation (discussed in detail in 3.3) that blurs the boundary between the virtual and physical interaction [29]. On the other hand, the aspects of mental and physical load of interactions in VR are not fully examined, and concerns of cybersickness should be addressed through further research [30].
3.3 Teleoperation
@@ 4 DISCUSSION
4.1 Mitigating Risks
The visualization of motion intention and object manipulation gives the operator a better understanding of the workflow. In the case of VR systems, one of the main advantages is the elimination of the need for physical expert presence [45]. In general, studies show that XR is a unique tool that allows the conduct of teleoperation, robot programming, and various operator-supporting functions in a safe and controlled manner. Although XR offers the possibility of finding a balance between operator safety and robot efficiency, the impact on physical and mental health from working with head-mounted displays in HRI tasks necessitates further studies. As mentioned in [30], there are currently no optimal solutions to address all possible side effects like muscle fatigue, motion sickness, and mental overload. Similarly, for human-robot collaborative tasks, the psychological factor remains a significant area of research. This includes the effort to cultivate trust between the agents, as well as methods to generate motion and trajectories that closely mimic human behavior [46].
4.2 Immersiveness
One of the bases for introducing XR is its immersive potential. The immersive environment allows users to perceive the spatial aspects of a robot and its surroundings more effectively. By providing a realistic and engaging experience, users can interact with and understand the robot’s movements and actions in a way that is not possible through traditional interfaces. This level of interaction and understanding can facilitate better demonstration quality, which is one of the most decisive factors in the effectiveness of robot policy learning [38, 47].
Given the enhanced visualization and situational awareness provided by immersive capabilities, task coordination is another area that can be improved for better collaboration between agents. The application of XR in the context of multi-robot systems can provide easy-to-use, intuitive methods for commanding multiple agents synchronously [48].
4.3 User-oriented Concepts
@@
4.4 From Lab to Industry
One of the primary obstacles preventing the adaptation of extended reality for the human-robot collaboration scenarios is setup costs, requiring substantial initial investment. Additionally, there is a lack of unified consensus within industry on the XR integration strategy and interface development, which understandably stems from the differences in industry-specific requirements and individual products. A possible future area of research could involve investigating the feasibility of utilizing the same XR interface across multiple manipulators with the flexibility to easily customize the interface for new robot- or product-specific characteristics. Furthermore, the implementation of XR is usually performed in a bundle with other digital technologies, as outlined in 2.3. Therefore, it entails additional investments and resources, and in fact, the effectiveness of XR becomes conditional on the development and maintenance of other technologies. Also, it is noteworthy that most of the reviewed works perform experiments in the laboratory. The prospect of transferring the developed methodologies and interfaces to real-world scenarios and industrial settings remains yet to be explored.
5 CONCLUSIONS